let ck>0 for k=1,2,....,n.show that( c1+c2+....cn)/√n≤[ c1^2+c2^2+....+cn^2]^1/2≤c1+c2+...+cn


1
Expert's answer
2022-03-16T17:55:02-0400

"\frac{c_1+c_2+...+c_n}{\sqrt{n}}\leq\sqrt{c_1^2+c_2^2+...+c_n^2}\leq \\
\leq c_1+c_2+...+c_n\\
c_k>0, k=1,2,...,n\\
\frac{c_1+c_2+...+c_n}{\sqrt{n}}\leq\sqrt{c_1^2+c_2^2+...+c_n^2}\\
(\frac{c_1+c_2+...+c_n}{\sqrt{n}})^2\leq(\sqrt{c_1^2+c_2^2+...+c_n^2})^2\\
\frac{(c_1+c_2+...+c_n)^2}{n}\leq c_1^2+c_2^2+...+c_n^2\\
(c_1+c_2+...+c_n)^2\leq n(c_1^2+c_2^2+...+c_n^2)"

If "n=1, c_1^2\leq c_1^2" true.

Lets "n=k, (c_1+c_2+...+c_k)^2\leq k(c_1^2+c_2^2+...+c_k^2)\\" .

Proof if "n=k+1"

"(c_1+c_2+...+c_{k+1})^2\leq\\\leq (k+1)(c_1^2+c_2^2+...+c_{k+1}^2)\\
(c_1+c_2+...+c_{k+1})^2=\\=((c_1+c_2+...+c_k)+c_{k+1})^2=\\
=(c_1+c_2+...+c_k)^2+\\+2(c_1+c_2+...+c_k)c_{k+1}+c_{k+1}^2
\leq\\\leq k(c_1^2+c_2^2+...+c_{k}^2)+\\
+2(c_1+c_2+...+c_k)c_{k+1}+c_{k+1}^2=\\
=k(c_1^2+c_2^2+...+c_{k}^2)+\\
+2c_1c_{k+1}+2c_2c_{k+1}+...+2c_kc_{k+1}+c_{k+1}^2\leq"

"(a-b)^2\geq 0\\
a^2-2ab+b^2\geq0\\
2ab\leq a^2+b^2"

"\leq k(c_1^2+c_2^2+...+c_{k}^2)+\\
+c_1^2+c^2_{k+1}+c_2^2+c^2_{k+1}+...+c^2_k+c^2_{k+1}+c_{k+1}^2=\\
=(k+1)(c_1^2+c_2^2+...+c_{k}^2)+(k+1)c^2_{k+1}=\\
(k+1)(c_1^2+c_2^2+...+c_{k}^2+c^2_{k+1})"

True

"\sqrt{c_1^2+c_2^2+...+c_n^2}\leq c_1+c_2+...+c_n\\
c_1^2+c_2^2+...+c_n^2\leq (c_1+c_2+...+c_n)^2\\
c_1^2+c_2^2+...+c_n^2\leq c_1^2+c_2^2+...+c_n^2+\\
+2c_1c_2+...+2c_1c_n+...+2c_{n-1}c_n\\
c_k>0, k=1,2,...,n\\
0<2c_1c_2+...+2c_1c_n+...+2c_{n-1}c_n"

True